OpenStudy (anonymous):

if r(t) and r'(t) are differentiable functions of (t), then find the value of d/dt[2r(t) x r'(t)]

OpenStudy (anonymous):

r(t) r'(t) = v(t) r''(t) = v'(t) = a'(t) so : 2r(t) X r'(t) = (2ry(t)vz(t) - 2rz(t)vy(t))i + (2rz(t)vx(t) - 2rx(t)vz(t))j + (2rx(t)vy(t) - 2ry(t)vx(t))k take the derivative of it : (2vy(t)vz(t) + 2ry(t)az(t) - 2vz(t)vy(t) - 2rz(t)ay(t))i + (2vz(t)vx(t) + 2rz(t)ax(t) - 2vx(t)vz(t) - 2rx(t)az(t))j + (2vx(t)vy(t) + 2rx(t)ay(t) -2vy(t)vx(t) -2ry(t)ax(t))k = (2ry(t)az(t) - 2rz(t)ay(t))i + (2rz(t)ax(t)- 2rx(t)az(t))j + (2rx(t)ay(t) - 2ry(t)ax(t))k

OpenStudy (anonymous):

or in your notations : (2ry(t)r''z(t) - 2rz(t)r''y(t))i + (2rz(t)r''x(t)- 2rx(t)r''z(t))j + (2rx(t)r''y(t) - 2ry(t)r''x(t))k

OpenStudy (anonymous):

please see the attachment and tell me the solution.

OpenStudy (jamesj):

The cross-product follows the chain rule; that is, let a(t) and b(t) be vector fields. Then \[\frac{d \ }{dt} [ a(t) \times b(t) ] = \frac{da}{dt} \times b + a \times \frac{db}{dt}\] Using this result, your problem should be straight forward.